OpenStudy (anonymous):

How many grams of water can be produced when 11.7 moles of ethane (C2H6) react with excess oxygen gas? Unbalanced equation: C2H6 + O2 → CO2 + H2O

OpenStudy (matt101):

Balanced: 2C2H6 + 7O2 → 4CO2 + 6H2O x/11.7 = 6/2 x = 35.1 moles of of water are produced 35.1 mol * 18 g/mol = 631.8 g of water are produced.

OpenStudy (anonymous):

where did you get 6/2 from? and what is x? where'd yu get 35.1? I am confused, I'm sorry.

OpenStudy (chmvijay):

see 2 moles of ethane gives =6 moles of water 11.7 moles of ethane gives = X moles of water (assume) cross multiplication =(6*11.7)/2 =35.1 moles

OpenStudy (matt101):

^ Sorry I wasn't more clear. Does that make sense now @andreadesirepen?